OpenStudy (anonymous):

Consider the equations y = |x − 1| and y = 3x + 2. The approximate solution of this system of equations is .

OpenStudy (irishboy123):

try solving these \(x − 1 = 3x + 2\) for \(x \ge 1\) and \(-(x- 1) = 3x + 2\) for \(x \lt 1\)